# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 4 Atomic Functions for NPDo
Armed with the general theoretical framework for the NPDo approach in section 3, in this section, we introduce the notion of atomic functions for NPDo, which serves as a singleton unit of function on $\mathbb{O}^{n\times k}_{\delta}$ for which the NPDo Ansatz holds and thus the NPDo approach is guaranteed to work, and more importantly, the NPDo approach works on any convex composition of atomic functions.
In what follows, we will first formulate two conditions that define atomic functions and explain why the NPDo approach will work, and then we give concrete examples of atomic functions that encompass nearly all practical ones that are in use today. We leave the investigation of how the NPDo approach works on convex compositions of atomic functions to section 5, along with a few convex compositions of our concrete atomic functions to guide the use of the general result.
@@ 4.1 Conditions on Atomic Functions
We are interested in functions $f$ defined on some neighborhood $\mathbb{O}^{n\times k}_{\delta}$ that satisfy
$\operatorname{tr}\left(P^{\operatorname{T}}\frac{\partial f(P)}{\partial P}\right)$ $=\gamma\,f(P)\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (4.1a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\alpha>0,\,\beta\geq 0$, and $\gamma=\alpha+\beta$ are constants that are dependent of $f$. Some subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$ is also involved and, as we commented in Remark 3.1(i) for the NPDo Ansatz, $\mathbb{P}$ should be sufficiently inclusive to serve the purpose of solving (1.1) with the function as objective, and for the case of $\mathbb{P}=\mathbb{O}^{n\times k}$, $Q$ can be taken to be $I_{k}$. More comments on this are in Remark 4.1(ii) below.
@@
A function $f$ defined on some neighborhood $\mathbb{O}^{n\times k}_{\delta}$ of $\mathbb{O}^{n\times k}$ is an atomic function for NPDo if there are constants $\alpha>0,\,\beta\geq 0$, and $\gamma=\alpha+\beta$ such that both conditions in (4.1) hold.
@@
Throughout this paper, we actually define two types of atomic functions. One type is what we just defined. It is for the NPDo approach. The other type will come in Part II later for the NEPv approach.
@@
    Theoretically, each of the two conditions in (4.1) has interest of its own. For example, equation (4.1a) is a partial differential equation (PDE) in its own right. In that regard, does it have a close form solution, given $\gamma\in\mathbb{R}$ (that is not necessarily nonnegative)? In this paper, we group the two together because later we need both to show that together they imply the NPDo Ansatz for an atomic function and thus the NPDo approach works. Also importantly, we need $\gamma=\alpha+\beta$.
@@
    How inclusive should $\mathbb{P}$ as a subset of $\mathbb{O}^{n\times k}$ be? Often certain necessary conditions for the maximizers of (1.1) with given atomic function as objective can be derived to limit the extent of searching. For example, as has been extensively exploited in [77, 67, 66, 74, 46], any maximizer $P_{*}$ must satisfy $P_{*}^{\operatorname{T}}D\succeq 0$ in the case where $f(P)$ contains $\operatorname{tr}(P^{\operatorname{T}}D)$ and increases as $\operatorname{tr}(P^{\operatorname{T}}D)$ does. In such a case, searching a maximizer can be naturally limited among those $P\in\mathbb{O}^{n\times k}$ such that $P^{\operatorname{T}}D\succeq 0$, i.e., $P\in\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. As a result, it suffices to just require that the equality and inequality in (4.1) hold for all $P,\,\widetilde{P}\in\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$. In our later concrete examples in subsection 4.2, equation (4.1a) even holds for all $P\in\mathbb{R}^{n\times k}$ for some atomic functions.
@@
$\operatorname{tr}\left(P^{\operatorname{T}}\frac{\partial g(P)}{\partial P}\right)$ $=s\gamma\,g(P)\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (4.2a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
yielding (4.2a). On the other hand, for $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, we have
@@
Given $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, suppose that (3.1) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}(P))\geq\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))+\eta$. We have by (4.1)
@@ 4.2 Concrete Atomic Functions
satisfy (4.1) and hence are atomic functions for NPDo. Therefore, by Theorem 4.2, the NPDo Ansatz holds for them. We point that the results we will prove in this subsection are actually for more general $P$, $\widehat{P}$ and $\widetilde{P}$ than required in Definition 4.1.
@@
      For $m>1$, if $P^{\operatorname{T}}D\succeq 0$, then
@@
      where $\widetilde{P}=\widehat{P}Q$ for $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}^{\operatorname{T}}D\succeq 0$.
In particular, the conditions in (4.1) hold with $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$, $\alpha=1$ and $\beta=m-1$, and thus $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ is an atomic function for NPDo. Furthermore just for $m=1$, (4.1b) becomes an identity and for the case we can also take $\mathbb{P}=\mathbb{O}^{n\times k}$ and $\widetilde{P}=\widehat{P}$ in (4.1).
@@
Now, we prove item (b). Equation (4.7) is easily verified. In general, for $m>1$, noticing the assumption $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$ for the case, we have, by Lemma B.6,
@@
In obtaining (4.12) by Lemma B.6, it is needed that $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$. This explains the necessity of having a strict subset $\mathbb{P}$ of $\mathbb{O}^{n\times k}$ and aligning $\widehat{P}$ to $\widetilde{P}\in\mathbb{P}$ by some $Q$ in (4.1) for defining atomic function for NPDo in general. For the purpose of maximizing $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$, given $P\in\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$ being the current approximation, to compute the next and hopefully improved approximation, the NPDo approach will seek $\widehat{P}$ to maximize $\operatorname{tr}\left(X^{\operatorname{T}}\frac{\partial\operatorname{tr}((P^{\operatorname{T}}D)^{m})}{\partial P}\right)$, or equivalently, $\operatorname{tr}(X^{\operatorname{T}}D(P^{\operatorname{T}}D)^{m-1})$, over $X\in\mathbb{O}^{n\times k}$, and hence $\widehat{P}$ is taken to be an orthogonal polar factor of $D(P^{\operatorname{T}}D)^{m-1}$. For that $\widehat{P}$, likely $\widehat{P}^{\operatorname{T}}D\not\succeq 0$, and hence necessarily $\widehat{P}$ needs to be aligned to $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ so that $\widetilde{P}^{\operatorname{T}}D\succeq 0$.
@@
    Let $P\in\mathbb{R}_{D+}^{n\times k}$, $\widehat{P}\in\mathbb{R}^{n\times k}$, and let $\widetilde{P}=\widehat{P}Q$, where $Q\in\mathbb{O}^{k\times}$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. We have $\widetilde{P}\in\mathbb{R}_{D+}^{n\times k}$ and
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}=\mathbb{O}_{D+}^{n\times k}$, $\alpha=1$ and $\beta=sm-1$, and thus $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ for $s>1$ is an atomic function for NPDo.
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}=\mathbb{O}^{n\times k}$, $\widetilde{P}=\widehat{P}$ , $\alpha=1$ and $\beta=2m-1$, and thus $\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ is an atomic function for NPDo.
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}=\mathbb{O}^{n\times k}$, $\alpha=1$ and $\beta=2sm-1$, and thus $[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ for $s>1$ is an atomic function for NPDo.
@@ 5 Convex Composition
The concrete atomic functions for NPDo in (4.5) provides a limited scope of objective functions for which the NPDo approach provably works. In this section, we will vastly expand the scope to include any convex composition of atomic functions, provided that some of the partial derivatives of the composing convex function are nonnegative.
Specifically, we are interested in a special case of optimization problem (1.1) on the Stiefel manifold $\mathbb{O}^{n\times k}$ where $f$ is a convex composition of atomic functions for NPDo, namely,
$\max_{P\in\mathbb{O}^{n\times k}}\,f(P)\quad\mbox{with}\quad f(P):=(\phi\circ T)(P)\equiv\phi(T(P)),$ (5.1)
where $T\,:\,P\in\mathbb{O}^{n\times k}\to T(P)\in\mathbb{R}^{N}$ whose components are atomic functions dependent of just a few or all columns of $P$, and $\phi\,:\,\mathbb{R}^{N}\to\mathbb{R}$ is convex and differentiable with partial derivatives, with respect to $\boldsymbol{x}=[x_{1},x_{2},\ldots,x_{N}]^{\operatorname{T}}\in\mathbb{R}^{N}$, denoted by
@@
$\mbox{with}\,\,\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\,\,P\in\mathbb{O}^{n\times k}.$ (5.4b)
@@
$\operatorname{tr}\left(P_{i}^{\operatorname{T}}\frac{\partial f_{i}(P_{i})}{\partial P_{i}}\right)$ $=\gamma_{i}\,f_{i}(P_{i})\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (5.5a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\alpha>0,\,\beta_{i}\geq 0$, and $\gamma_{i}=\alpha+\beta_{i}$ are constants, $\widehat{P}_{i}$ and $\widetilde{P}_{i}$ are the submatrices of $\widehat{P}$ and $\widetilde{P}$, respectively, with the same column indices taken as $P_{i}$ to $P$. It is important to keep in mind that some of the inequalities in (5.5b) may actually be equalities, e.g., for $f_{i}(P_{i})=\operatorname{tr}(P_{i}^{\operatorname{T}}D_{i})$ it is an equality by Theorem 4.3.
@@
Given $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, suppose that (3.1) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}(P))\geq\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))+\eta$. Write
@@
$T_{1}\,:\,P\in\mathbb{O}^{n\times k}\to T_{1}(P):=\begin{bmatrix}\operatorname{tr}(P_{1}^{\operatorname{T}}A_{1}P_{1})\\ \vdots\\ \operatorname{tr}(P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell})\\ \operatorname{tr}(P_{\ell+1}^{\operatorname{T}}D_{1})\\ \vdots\\ \operatorname{tr}(P_{\ell+t}^{\operatorname{T}}D_{t})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.6)
Either $\ell=0$ or $t=0$ (i.e., TrCP in Table 1) is allowed. If all $A_{i}\succeq 0$, then the consistency conditions in (5.5) are satisfied with $\mathbb{P}=\mathbb{O}^{n\times k}$, $Q=I_{k}$, $\alpha=1$, $\beta_{i}=1$ for $1\leq i\leq\ell$ and $\beta_{\ell+j}=0$ for $1\leq j\leq t$, by Theorems 4.3 and 4.4. In particular, now (5.5b) for $\ell+1\leq i\leq\ell+t$ are equalities. Thus Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$. Two existing special cases of $T_{1}$ are
@@
Despite that we can take $\mathbb{P}=\mathbb{O}^{n\times k}$ here, with favorable compositions of $P_{i}$ as submatrices of $P$ in (5.6), we can find a better $Q$, other than $I_{k}$, so that the objective value increases more than the NPDo Ansatz suggests. Here are two of them:
@@
Now determine $Q$ implicitly by taking just one index $j$ from each $\mathbb{I}_{q}$ for $1\leq q\leq\tau$ and letting $\widetilde{P}_{\ell+j}=\widehat{P}_{\ell+j}S_{j}$ where $S_{j}$ is an orthogonal polar factor of
@@
$T_{2}\,:\,P\in\mathbb{O}^{n\times k}\to T_{2}(P):=\begin{bmatrix}\|P_{1}^{\operatorname{T}}A_{1}P_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell}\|_{\operatorname{F}}^{2}\\ \|P_{\ell+1}^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P_{\ell+t}^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.10)
@@
Hence if all $A_{i}\succeq 0$, then the consistency conditions in (5.5) are satisfied with $\mathbb{P}=\mathbb{O}^{n\times k}$, $Q=I_{k}$, $\alpha=1$, $\beta_{i}=3$ for $1\leq i\leq\ell$, and $\beta_{\ell+j}=1$ for $1\leq j\leq t$. Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell+t$.

### v4

@@ 4 Atomic Functions for NPDo
Armed with the general theoretical framework for the NPDo approach in section 3, in this section, we introduce the notion of atomic functions for NPDo, which serves as a singleton unit of function on ${\rm St}_{\delta}(k,n)$ for which the NPDo Ansatz holds and thus the NPDo approach is guaranteed to work for solving (1.1), and more importantly, the NPDo approach works on any convex composition of atomic functions.
In what follows, we will first formulate two conditions that define atomic functions and explain why the NPDo approach will work on the atomic functions, and then we give concrete examples of atomic functions that encompass nearly all practical ones that are in use today. We leave the investigation of how the NPDo approach works on convex compositions of atomic functions to section 5, along with a few convex compositions of our concrete atomic functions to guide the use of the general result.
@@ 4.1 Conditions on Atomic Functions
We are interested in functions $f$ defined on some neighborhood ${\rm St}_{\delta}(k,n)$ of the Stiefel manifold ${\rm St}(k,n)$ that satisfy
$\operatorname{tr}\left(P^{\operatorname{T}}\frac{\partial f(P)}{\partial P}\right)$ $=\gamma\,f(P)\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (4.1a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\alpha>0,\,\beta\geq 0$, and $\gamma=\alpha+\beta$ are constants that are dependent of $f$. Some subset $\mathbb{P}$ of ${\rm St}(k,n)$ is also involved and, as we commented in Remark 3.1(i) for the NPDo Ansatz, $\mathbb{P}$ should be sufficiently inclusive to serve the purpose of solving (1.1) with the function as objective, and for the case of $\mathbb{P}={\rm St}(k,n)$, $Q$ can be taken to be $I_{k}$. More comments on this are in Remark 4.1(ii) below.
@@
A function $f$ defined on some neighborhood ${\rm St}_{\delta}(k,n)$ of ${\rm St}(k,n)$ is an atomic function for NPDo if there are constants $\alpha>0,\,\beta\geq 0$, and $\gamma=\alpha+\beta$ such that both conditions in (4.1) hold.
@@
Throughout this paper, we actually define two types of atomic functions. One type is what we just defined in Definition 4.1. It is for the NPDo approach. The other type will come in Part II later for the NEPv approach.
@@
    Theoretically, each of the two conditions in (4.1) has interest of its own. For example, equation (4.1a) is a partial differential equation (PDE) in its own right. In that regard, a natural question arises: does it have a close form solution, given $\gamma\in\mathbb{R}$ (that is not necessarily nonnegative)? In this paper, we group the two together because later we need both to show that together they imply the NPDo Ansatz for an atomic function and thus the NPDo approach works. Also importantly, we need $\gamma=\alpha+\beta$.
@@
    How inclusive should $\mathbb{P}$ as a subset of ${\rm St}(k,n)$ be? Often certain necessary conditions for the maximizers of (1.1) with given atomic function as objective can be derived to limit the extent of searching. For example, as has been extensively exploited in [77, 67, 66, 74, 46], any maximizer $P_{*}$ must satisfy $P_{*}^{\operatorname{T}}D\succeq 0$ in the case where $f(P)$ contains $\operatorname{tr}(P^{\operatorname{T}}D)$ and increases as $\operatorname{tr}(P^{\operatorname{T}}D)$ does. In such a case, searching a maximizer can be naturally limited among those $P\in{\rm St}(k,n)$ such that $P^{\operatorname{T}}D\succeq 0$, i.e., $P\in\mathbb{P}={\rm St}(k,n)_{D+}$. As a result, it suffices to just require that the equality and inequality in (4.1) hold for all $P,\,\widetilde{P}\in\mathbb{P}={\rm St}(k,n)_{D+}$. In our later concrete examples in subsection 4.2, equation (4.1a) even holds for all $P\in\mathbb{R}^{n\times k}$ for some atomic functions.
@@
$\operatorname{tr}\left(P^{\operatorname{T}}\frac{\partial g(P)}{\partial P}\right)$ $=s\gamma\,g(P)\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (4.2a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
yielding (4.2a). On the other hand, for $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, we have
@@
Given $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and $\widehat{P}\in{\rm St}(k,n)$, suppose that (3.1) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}(P))\geq\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))+\eta$. We have by (4.1)
@@ 4.2 Concrete Atomic Functions
satisfy (4.1) and hence are atomic functions for NPDo. Therefore, by Theorem 4.2, the NPDo Ansatz holds for them. We point out that the results we will prove in this subsection are actually for more general $P$, $\widehat{P}$ and $\widetilde{P}$ than required in Definition 4.1.
@@
      For $m\geq 1$, if $P^{\operatorname{T}}D\succeq 0$, then
@@
      where $\widetilde{P}=\widehat{P}Q$ for $Q\in{\rm St}(k,k)$ such that $\widetilde{P}^{\operatorname{T}}D\succeq 0$.
In particular, the conditions in (4.1) hold with $\mathbb{P}={\rm St}(k,n)_{D+}$, $\alpha=1$ and $\beta=m-1$, and thus $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$ is an atomic function for NPDo.
Just for the case $m=1$, we can also take $\mathbb{P}={\rm St}(k,n)$ and $\widetilde{P}=\widehat{P}$ in (4.1), and then (4.1b) becomes an equality. Therefore, $\operatorname{tr}(P^{\operatorname{T}}D)$ is also an atomic function for NPDo with $\mathbb{P}={\rm St}(k,n)$ and $Q=I_{k}$ in the definition.
@@
Now, we prove item (b). Equation (4.7) which is for $m=1$ is easily verified. In general, for $m>1$, noticing the assumption $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$ for the case, we have, by Lemma B.6,
@@
In obtaining (4.12) by Lemma B.6, it is needed that $P^{\operatorname{T}}D\succeq 0$ and $\widetilde{P}^{\operatorname{T}}D=Q^{\operatorname{T}}(\widehat{P}^{\operatorname{T}}D)\succeq 0$. This explains the necessity of having a strict subset $\mathbb{P}$ of ${\rm St}(k,n)$ and aligning $\widehat{P}$ to $\widetilde{P}\in\mathbb{P}$ by some $Q$ in (4.1) for defining atomic function for NPDo in general. For the purpose of maximizing $\operatorname{tr}((P^{\operatorname{T}}D)^{m})$, given $P\in\mathbb{P}={\rm St}(k,n)_{D+}$ being the current approximation, to compute the next and hopefully improved approximation, the NPDo approach will seek $\widehat{P}$ to maximize $\operatorname{tr}\left(X^{\operatorname{T}}\frac{\partial\operatorname{tr}((P^{\operatorname{T}}D)^{m})}{\partial P}\right)$, or equivalently, $\operatorname{tr}(X^{\operatorname{T}}D(P^{\operatorname{T}}D)^{m-1})$, over $X\in{\rm St}(k,n)$, and hence $\widehat{P}$ is taken to be an orthogonal polar factor of $D(P^{\operatorname{T}}D)^{m-1}$. For that $\widehat{P}$, likely $\widehat{P}^{\operatorname{T}}D\not\succeq 0$, and hence necessarily $\widehat{P}$ needs to be aligned to $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ so that $\widetilde{P}^{\operatorname{T}}D\succeq 0$.
@@
    Let $P\in\mathbb{R}_{D+}^{n\times k}$, $\widehat{P}\in\mathbb{R}^{n\times k}$, and let $\widetilde{P}=\widehat{P}Q$, where $Q\in{\rm St}(k,k)$ is an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}D$. We have $\widetilde{P}\in\mathbb{R}_{D+}^{n\times k}$ and
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}={\rm St}(k,n)_{D+}$, $\alpha=1$ and $\beta=sm-1$, and thus $[\operatorname{tr}((P^{\operatorname{T}}D)^{m})]^{s}$ for $s>1$ is an atomic function for NPDo.
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$ and $\widetilde{P}=\widehat{P}$ , $\alpha=1$ and $\beta=2m-1$, and thus $\operatorname{tr}((P^{\operatorname{T}}AP)^{m})$ is an atomic function for NPDo.
@@
In particular, the conditions in (4.1) hold with $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$, $\alpha=1$ and $\beta=2sm-1$, and thus $[\operatorname{tr}((P^{\operatorname{T}}AP)^{m})]^{s}$ for $s>1$ is an atomic function for NPDo.
@@ 5 Convex Composition
The concrete atomic functions for NPDo in (4.5) provides a limited collection of objective functions for which the NPDo approach provably works. In this section, we will vastly expand the collection to include any convex composition of atomic functions, provided that some of the partial derivatives of the composing convex function are nonnegative.
Specifically, we are interested in a special case of optimization problem (1.1) on the Stiefel manifold ${\rm St}(k,n)$ where $f$ is a convex composition of atomic functions for NPDo, namely,
$\max_{P\in{\rm St}(k,n)}\,f(P)\quad\mbox{with}\quad f(P):=(\phi\circ T)(P)\equiv\phi(T(P)),$ (5.1)
where $T\,:\,P\in{\rm St}(k,n)\to T(P)\in\mathfrak{D}\subseteq\mathbb{R}^{N}$ whose components are atomic functions dependent of just a few or all columns of $P$, and $\phi\,:\,\mathfrak{D}\to\mathbb{R}$ is convex and differentiable. Denote the partial derivatives of $\phi$ with respect to $\boldsymbol{x}=[x_{1},x_{2},\ldots,x_{N}]^{\operatorname{T}}\in\mathfrak{D}\subseteq\mathbb{R}^{N}$ by
@@
$\mbox{with}\,\,\Lambda^{\operatorname{T}}=\Lambda\in\mathbb{R}^{k\times k},\,\,P\in{\rm St}(k,n).$ (5.4b)
@@
$\operatorname{tr}\left(P_{i}^{\operatorname{T}}\frac{\partial f_{i}(P_{i})}{\partial P_{i}}\right)$ $=\gamma_{i}\,f_{i}(P_{i})\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (5.5a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\alpha>0,\,\beta_{i}\geq 0$, and $\gamma_{i}=\alpha+\beta_{i}$ are constants, $\widehat{P}_{i}$ and $\widetilde{P}_{i}$ are the submatrices of $\widehat{P}$ and $\widetilde{P}$, respectively, with the same column indices as $P_{i}$ to $P$. It is important to keep in mind that some of the inequalities in (5.5b) may actually be equalities, e.g., for $f_{i}(P_{i})=\operatorname{tr}(P_{i}^{\operatorname{T}}D_{i})$ it is an equality by Theorem 4.3.
@@
Given $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and $\widehat{P}\in{\rm St}(k,n)$, suppose that (3.1) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}\mathscr{H}(P))\geq\operatorname{tr}(P^{\operatorname{T}}\mathscr{H}(P))+\eta$. Write
@@
$T_{1}\,:\,P\in{\rm St}(k,n)\to T_{1}(P):=\begin{bmatrix}\operatorname{tr}(P_{1}^{\operatorname{T}}A_{1}P_{1})\\ \vdots\\ \operatorname{tr}(P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell})\\ \operatorname{tr}(P_{\ell+1}^{\operatorname{T}}D_{1})\\ \vdots\\ \operatorname{tr}(P_{\ell+t}^{\operatorname{T}}D_{t})\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.6)
Either $\ell=0$ or $t=0$ (i.e., TrCP in Table 1) is allowed. If all $A_{i}\succeq 0$, then the consistency conditions in (5.5) are satisfied with $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$, $\alpha=1$, $\beta_{i}=1$ for $1\leq i\leq\ell$ and $\beta_{\ell+j}=0$ for $1\leq j\leq t$, by Theorems 4.3 and 4.4. In particular, now (5.5b) for $\ell+1\leq i\leq\ell+t$ are equalities. Thus Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell$. Two existing special cases of $T_{1}$ are
@@
Despite that we can take $\mathbb{P}={\rm St}(k,n)$ and $Q=I_{k}$ here, with favorable compositions of $P_{i}$ as submatrices of $P$, we can find a better $Q$, other than $I_{k}$, so that the objective value increases more than the NPDo Ansatz suggests. Here are two of them:
@@
Now determine $Q$ implicitly by taking just one index $j$ from each $\mathbb{I}_{q}$ for $1\leq q\leq\tau$ and letting $\widetilde{P}_{\ell+j}=\widehat{P}_{\ell+j}S_{q}$ where $S_{q}$ is an orthogonal polar factor of
@@
$T_{2}\,:\,P\in{\rm St}(k,n)\to T_{2}(P):=\begin{bmatrix}\|P_{1}^{\operatorname{T}}A_{1}P_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P_{\ell}^{\operatorname{T}}A_{\ell}P_{\ell}\|_{\operatorname{F}}^{2}\\ \|P_{\ell+1}^{\operatorname{T}}D_{1}\|_{\operatorname{F}}^{2}\\ \vdots\\ \|P_{\ell+t}^{\operatorname{T}}D_{t}\|_{\operatorname{F}}^{2}\\ \end{bmatrix}\in\mathbb{R}^{\ell+t}.$ (5.10)
@@
Hence if all $A_{i}\succeq 0$, then the consistency conditions in (5.5) are satisfied with $\mathbb{P}={\rm St}(k,n)$, $Q=I_{k}$, $\alpha=1$, $\beta_{i}=3$ for $1\leq i\leq\ell$, and $\beta_{\ell+j}=1$ for $1\leq j\leq t$. Theorem 5.1 applies, assuming $\phi_{j}(\boldsymbol{x})\geq 0$ for $1\leq j\leq\ell+t$.